# Asymptotic condition numbers for linear ODEs

### v1

@@ Abstract.
We are interested in the (relative) conditioning of the linear problem $y_{0}\mapsto\mathrm{e}^{tA}y_{0}$, i.e. the conditioning of the action of the matrix exponential $\mathrm{e}^{tA}$ on a vector with respect to perturbations of this vector. The present paper is a qualitative study of the long-time behavior of this conditioning. In other words, we are interested to study the propagation to the solution $y(t)$ of perturbations of the initial value for a linear ODE $y^{\prime}(t)=Ay(t)$, by measuring these perturbations with relative errors instead of absolute errors. We introduce three condition numbers: the first considers a specific initial value and a specific direction of perturbation; the second considers a specific initial value and the worst case by varying the direction of perturbation; and the third considers the worst case by varying both the initial value and the direction of perturbation. The long-time behaviors of these three condition numbers are studied.
@@ 1. Introduction
We are interested to understand how a perturbation of the initial value $y_{0}$ of the linear $n-$dimensional Ordinary Differential Equation (ODE)
@@
where $A\in\mathbb{C}^{n\times n}$ and $y_{0},y(t)\in\mathbb{C}^{n}$, is propagated to the solution $y(t)=\mathrm{e}^{tA}y_{0}$ of (1) over a long time $t$. This perturbation, propagating along the solution, is measured by a relative error rather than an absolute error. In other words, we study the (relative) conditioning of the linear problem
$y_{0}\mapsto\mathrm{e}^{tA}y_{0}$ (2)
@@
The conditioning of the problem (2) has received little attention in literature despite of its significance in the context of ODEs. In contrast, the study of the conditioning of the matrix exponential $\mathrm{e}^{tA}$ is quite extensive; see, for example, [10], [15], [9], [11], [14], [2], [16], [3], [5], and [1]. However, none of these papers examined how the conditioning depends on $t$, a crucial aspect for ODEs. This type of analysis, for $A$ normal, was addressed in [12] for the problem (2) and in [7] for the problem $A\mapsto\mathrm{e}^{tA}y_{0}$. In the present paper, for the problem (2), we consider a general matrix $A$.
The present paper complements the work presented in [13]. The paper [13] offers a non-technical exposition without proofs of the main results found in the present paper. It includes motivations, practical examples, and numerical tests. Consequently, these aspects are not addressed in the present paper, which focuses on providing the proofs of all results and the necessary theoretical development for obtaining them.
@@ 1.1. The condition numbers
[…]
$\varepsilon=\frac{\left\|\widetilde{y}_{0}-y_{0}\right\|}{\left\|y_{0}\right\|}$
@@
$\delta\left(t\right)=\frac{\left\|\widetilde{y}\left(t\right)-y\left(t\right)\right\|}{\left\|y\left(t\right)\right\|}$
@@
where $\widehat{z}_{0}\in\mathbb{C}^{n}$ with $\|\widehat{z}_{0}\|=1$ is the direction of perturbation, we obtain
@@
with $\widehat{y}_{0}:=\frac{y_{0}}{\left\|y_{0}\right\|}$, is called the directional condition number of the problem (2).
Along with the condition number (3), we also introduce other two condition numbers: the condition number of the problem (2) (see [4] for the definition of condition number of a general problem) given by
@@
[…]
@@
Observe that $K\left(t,y_{0}\right)$ is the worst condition number $K(t,y_{0},\widehat{z}_{0})$ by varying $\widehat{z}_{0}$ and $K(t)$ is the worst condition number $K(t,y_{0})$ by varying $y_{0}$, i.e. the worst condition number $K(t,y_{0},\widehat{z}_{0})$ by varying both $y_{0}$ and $\widehat{z}_{0}$.
@@
The paper [12] studied the condition numbers (3), (4) and (5) in case of $A$ normal and the Euclidean norm as vector norm on $\mathbb{R}^{n}$. The present paper study the general case where $A$ (complex) and the vector norm are arbitrary.

### v2

# Asymptotic condition numbers for linear ordinary differential equations

@@ Abstract.
We are interested in the relative conditioning of the problem $y_{0}\mapsto\mathrm{e}^{tA}y_{0}$, i.e., the relative conditioning of the action of the matrix exponential $\mathrm{e}^{tA}$ on a vector with respect to perturbations of this vector. The present paper is a qualitative study of the long-time behavior of this conditioning. In other words, we are interested in studying the propagation to the solution $y(t)$ of perturbations of the initial value for a linear ordinary differential equation $y^{\prime}(t)=Ay(t)$, by measuring these perturbations with relative errors. We introduce three condition numbers: the first considers a specific initial value and a specific direction of perturbation; the second considers a specific initial value and the worst case by varying the direction of perturbation; and the third considers the worst case by varying both the initial value and the direction of perturbation. The long-time behaviors of these three condition numbers are studied.
@@ 1. Introduction
We are interested in understanding how a perturbation of the initial value $y_{0}$ of the linear $n$-dimensional Ordinary Differential Equation (ODE)
@@
where $A\in\mathbb{C}^{n\times n}$ and $y_{0},y(t)\in\mathbb{C}^{n}$, is propagated to the solution $y(t)=\mathrm{e}^{tA}y_{0}$ of (1) over a long time interval. This perturbation, propagating along the solution, is measured by a relative error. In other words, we study the relative conditioning of the problem
$y_{0}\mapsto y(t)=\mathrm{e}^{tA}y_{0}$ (2)
@@
The relative conditioning of the matrix exponential function, i.e., the relative conditioning of the problem $A\mapsto\mathrm{e}^{A}$, or the problem
$A\mapsto\mathrm{e}^{tA}$ (3)
involving the time $t$, has been extensively studied: see [9], [15], [8], [10], [14], [2], [16], [3], [5], and [1]. In the study of time evolutions, an important aspect is to understand how the relative conditioning depends on $t$. Many of the papers cited above have examined this issue for the problem (3). For a normal matrix $A$, it is known that the relative condition number of (3) grows linearly with $t$. On the other hand, for a general matrix $A$, the exact order of growth with $t$ and its dependence on the matrix $A$ are not known: we have only polynomial lower bounds and exponential upper bounds in $t$ for the relative condition number.
It is also of interest to study the relative conditioning of the action of the matrix exponential on a vector. This is particularly important in the context of ODEs (1), where the solution is given by the action of the matrix exponentials $\mathrm{e}^{tA}$ on a known initial value $y_{0}$. In this context, we can consider the relative conditioning of the problems (2) and
$A\mapsto\mathrm{e}^{tA}y_{0},$ (4)
which are more important than the problem (3), since $y_{0}$ is involved. Despite the importance, there has been little attention in the literature to the relative conditioning of these problems.
For the case of $A$ normal, an analysis of the relative conditioning, focused on time $t$, was carried out in [6] for the problem (4).
The relative conditioning of problem (2) can be perceived, at first glance, as a trivial issue: (2) is a linear problem and its condition number can be immediately determined and computed. This perception is especially reinforced when the relative conditioning of (2) is compared with the relative conditioning of the non-linear problems (3) and (4). However, the relative conditioning of (2) ceases to be trivial once the time $t$ is taken into account, and we want to understand how it depends on $t$.
Analyzing how the relative conditioning of (2) depends on $t$ can fill a gap in our understanding of linear dynamics. In fact, while it is well understood how the absolute conditioning of (2) depends on $t$, i.e., how absolute errors due to perturbations of $y_{0}$ propagate to $y(t)$ (they are governed by the real parts of the rightmost eigenvalues of $A$ for large $t$, and by the pseudospectra of $A$ for non-large $t$), how the relative conditioning of (2) depends on $t$, i.e., how relative errors propagate, is far less understood, even for large $t$. Such an analysis was carried out in [11] for the simple case of $A$ normal. In the present paper, we carry out this analysis for a general ODE (1), by considering the relative conditioning for large $t$. The analysis is far from trivial, as evidenced by the length of this paper and its sequels [12] and [13].
@@ 1.1. The condition numbers
Suppose that the initial value $y_{0}\neq 0$ in (1) is perturbed to $\widetilde{y}_{0}$ and, as a consequence, the solution $y$ is perturbed to $\widetilde{y}$. Let $\left\|\ \cdot\ \right\|$ be an arbitrary vector norm on $\mathbb{C}^{n}$. We introduce the normwise relative error
$\varepsilon:=\frac{\left\|\widetilde{y}_{0}-y_{0}\right\|}{\left\|y_{0}\right\|}$
@@
$\delta\left(t\right):=\frac{\left\|\widetilde{y}\left(t\right)-y\left(t\right)\right\|}{\left\|y\left(t\right)\right\|}$
@@
where $\widehat{z}_{0}\in\mathbb{C}^{n}$, with $\|\widehat{z}_{0}\|=1$, is the direction of perturbation, we obtain
@@
with $\widehat{y}_{0}:=\frac{y_{0}}{\left\|y_{0}\right\|}$ the normalized initial value. The number $K\left(t,y_{0},\widehat{z}_{0}\right)$ is called the directional pointwise condition number of the problem (2): it is called “directional” because it depends on $\widehat{z}_{0}$, and “pointwise” because it depends on $y_{0}$.
Along with the condition number (5), we also introduce two other condition numbers:
  • the pointwise condition number of the problem (2) given by
@@
    where $\left\|\mathrm{e}^{tA}\right\|$ is the matrix norm of $\mathrm{e}^{tA}$ induced by the vector norm $\left\|\ \cdot\ \right\|$ (see [4] for the definition of condition number of a general problem, which corresponds to the pointwise condition number);
  • the global condition number of the problem (2) given by
@@
Observe that $K\left(t,y_{0}\right)$ is the worst $K(t,y_{0},\widehat{z}_{0})$ by varying $\widehat{z}_{0}$, and $K(t)$ is the worst $K(t,y_{0})$ by varying $y_{0}$, i.e., the worst $K(t,y_{0},\widehat{z}_{0})$ by varying both $y_{0}$ and $\widehat{z}_{0}$.
@@
The paper [11] studied the condition numbers (5), (6) and (7) in the particular case of $A$ normal. The present paper studies the general case.
The aim of the present paper is to analyze the asymptotic (long-time) behavior of the three condition numbers $K(t,y_{0},\widehat{z}_{0})$, $K(t,y_{0})$ and $K(t)$, i.e., their behavior as $t$ approaches infinity (becomes large).
1.2. Plan of the paper
Besides this introduction, the paper contains five sections and three appendices.
Section 2 develops notions and notations for understanding the asymptotic forms of the three condition numbers. Section 3 analyzes the asymptotic behaviors of the condition numbers $K(t,y_{0},\widehat{z}_{0})$ and $K(t,y_{0})$, by introducing the asymptotic condition numbers $K_{\infty}(t,y_{0},\widehat{z}_{0})$ and $K_{\infty}(t,y_{0})$. Section 4 analyzes the asymptotic behavior of the condition number $K(t)$, by introducing the asymptotic condition numbers $K_{\infty}^{+}(t)$ and $K_{\infty}(t)$. Section 5 introduces the Rightmost Last Generalized Eigenvector (RLGE) condition and summarizes the most important results. Conclusions are in Section 6.
The three appendices contain the more technical material. They should be consulted as needed while reading the paper, and read in full only by readers interested in the mathematical details. Appendix A develops a suitable formula for the matrix exponential $\mathrm{e}^{tA}$ in the general non-diagonalizable case along with other fundamental material related to this formula. Appendix B investigates the properties of the key matrices $Q_{jl}(t)$ that determine the asymptotic behavior of the condition numbers. Appendix C analyzes the matrices $Q_{j}^{e}(t)$, which are important for defining the asymptotic condition number $K_{\infty}(t)$ related to $K(t)$.
The present paper has two sequels. The first is [12], which develops the results of this paper in depth, for a real ODE (1) in a generic case. The second is [13], which presents extensive experimental tests, applications to real-world systems, and other issues, such as the non-asymptotic behavior of the condition numbers and how rapidly the asymptotic behavior sets in, also in relation to the non-normality of the matrix $A$. As a consequence, we will not discuss such practical questions in the present paper nor in [12].
In any case, the present paper reaches a conclusion by fully characterizing the asymptotic behavior of the condition numbers (5), (6) and (7). Moreover, it contains in Sections 4 and 5 some numerical examples illustrating the results obtained in this paper. Papers [12] and [13] are further developments that cannot be included here for obvious space constraints.
In what follows, we often refer to a generic case for an element $v$ of a finite-dimensional space $V$. By this, we mean that $v$ satisfies a property which is not satisfied only on a manifold $M$ of $V$ with $\mathrm{dim}M<\mathrm{dim}V$. Equivalently, if $v$ is drawn at random from $V$ (with respect to any distribution absolutely continuous with Lebesgue measure), the generic case holds with probability $1$.
2. Asymptotic forms
In the next Sections 3 and 4, we analyze the asymptotic behavior of the three condition numbers $K(t,y_{0},\widehat{z}_{0})$, $K(t,y_{0})$ and $K(t)$. This analysis determines the asymptotic forms of $\mathrm{e}^{tA}\widehat{z}_{0}$ and $\mathrm{e}^{tA}\widehat{y}_{0}$ in (5), of $\mathrm{e}^{tA}$ and $\mathrm{e}^{tA}\widehat{y}_{0}$ in (6), and of $\mathrm{e}^{tA}$ and $\mathrm{e}^{-tA}$ in (7). Then, the asymptotic behavior of the condition numbers is described by inserting in (5), (6) and (7) these asymptotic forms and by defining as asymptotic condition numbers the new expressions obtained by these substitutions.
All of this might seem straightforward, but from a rigorous mathematical perspective it is not a simple task. Specifically, the following points should be remarked.
  • It is necessary to precisely define what we mean by asymptotic form and how to determine it. The determination of the asymptotic forms is complicated in the non-diagonalizable case, where the Jordan Canonical Form of $A$ and generalized eigenvectors are involved. Moreover, proving that a given candidate for an asymptotic form is indeed an asymptotic form in our definition requires a certain mathematical effort: see Remark 6 at the end of this section.
  • We also want to quantify how dominant the asymptotic forms are at finite times (these quantifications are used in the sequel papers [12] and [13]).
  • The core of our analysis is the qualitative study of the asymptotic condition numbers. Understanding them requires significant mathematical effort, in particular in the case of rightmost complex eigenvalues (see the sequel paper [12]).
  • If we are interested in defining asymptotic condition numbers for the problem (2), the asymptotic condition numbers derived from (6) and (7) as $t\rightarrow+\infty$ may not be appropriate, since they represent the asymptotic worst cases of (5), by varying $\widehat{z}_{0}$ only and both $\widehat{z}_{0}$ and $y_{0}$, respectively. Instead, we may be concerned with the worst asymptotic cases of (5). In other words: do “asymptotic” and “worst” commute?
In this section, we determine the asymptotic forms of $\mathrm{e}^{tA}$ and $\mathrm{e}^{tA}u$, where $u\in\mathbb{C}^{n}$. In order to define them, some preliminary notions and results need to be introduced.
2.1. Notations $\sim$ and $\approx$
In this subsection, we make precise what we mean by asymptotic form.
Let $f(t)$ and $g(t)$ be scalar, vector or matrix functions of $t\in\mathbb{R}$. For $g$ such that $g(t)\neq 0$ for $t$ in a neighborhood of $\infty$, we write
$f\left(t\right)\sim g\left(t\right),\ t\rightarrow+\infty,$ (8)
when
In case of scalars, vectors and matrices, $\|\;\cdot\;\|$ denotes, respectively, the modulus, a vector norm and a matrix norm.
We interpret (8) as indicating that $g$ is an asymptotic form of $f$.
when
Observe that (9) means that the relative error of $f$ with respect to $g$ at the time $t$ is not larger than $\epsilon$. The notation (9) serves to quantify how dominant the asymptotic form $g$ of $f$ is at the finite time $t$.